Mikaela1509:

2/5w = −10. Please help I don't understand this

snowflake0531:

is it \[\frac{2}{5}x = -10\] or \[\frac{2}{5w}=-10\]

snowflake0531:

w**

Mikaela1509:

\[\frac{ 2 }{ 5 }w = -10\]

Mikaela1509:

the thing we need to find is what w =

umm:

I'll hop on PC and help you in a second.

Mikaela1509:

ok

umm:

Sorry for the wait. Might take a bit..

Krissy525:

so it would be the first thing i said 2/5w=−10 multiply by 5/2 on each side then you should get w=-25

umm:

Given: \[\frac{ 2 }{ 5 }w=-10\] First thing you'll need to do is multiply both sides by 5/2, so like this:\[\left(\begin{matrix}5 \\ 2\end{matrix}\right)*\left(\begin{matrix}2 \\ 5w\end{matrix}\right)=\left(\begin{matrix}5 \\ 2\end{matrix}\right)*(-10)\] \[w=−25\]Therefore, your answer is \[w=−25\]

umm:

Not sure how to do the fraction in the equation form for that, my bad.

crispyrat:

so 2/5w=-10 you multiply both sides by 5/2 to get one w on the left side And thus you get w=-50/2 w=-25 :)

Mikaela1509:

thx